[Ncep.nhc.nco_contacts] HSU Shift Log -- Techincal Edition -- Tue May 24, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue May 24 19:36:51 UTC 2016


HSU Shift Log for Tue May 24 2016 15:36:21 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: UKMET data did not come in at first at 12z. Dave Z.  
brought it in manually and it seemed to work. He thought it had to do with  
a switch to a back-up server.

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160524/968dd3b3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list